what does a = in this
2a+5=a+9

Answer by texttutoring(324)

2a+5=a+9
Think of an algebraic equation like a balance or a teeter totter. Any operation you do to one side, you must do to the other side. Let's try to get a on the left side, and all the numbers on the right side:

2a+5=a+9
2a-a+5 = a - a + 9
a+5 = 0+9
a+5 = 9
a = 9 - 5
a = 4

Let's check it to see if it's right:
2a+5=a+9
2(4) + 5 = 4 + 9
8+5 = 4+ 9
13 = 13

Yes, it's correct. a=4

Hope that helps!